A box contains 6 red balls, 4 white balls and 5 blue balls. Determine the probability of chasing two red balls and 1 white ball if the balls are not replaced after being drawn using combination method


1
Expert's answer
2022-04-29T13:46:01-0400
"P(2R\&\ 1W)=\dfrac{\dbinom{6}{2}\dbinom{4}{1}\dbinom{5}{0}}{\dbinom{6+4+5}{2+1}}"

"=\dfrac{15(4)(1)}{455}=\dfrac{12}{91}"